Please solve the following equations by factors (Please do not use the quadratic formula)
x squared -2x-3=0

4x squared -29x+7=0

Answer by ewatrrr(24785)


Hi,
x^-2x - 3 = 0
factor and solve
x^-2x - 3 = 0 Note: factoring will have the format (x - ?)(x + ?)
(x-3)(x+1)= 0 this format works as -3x + x = -2x
(x-3)= 0
x = 3
(x+1)= 0
x = -1